Information on Bowgie Chemicals direct materials costs follows.

 





Actual quantities of direct materials used
30,800
Actual costs of direct materials used$137,800
Standard price per unit of direct materials$4.28
Flexible budget for direct materials$121,300

 

Bowgie Chemicals has no materials inventories.

 

Required:
a. 
Prepare a short report for management showing Bowgie Chemicals’s direct materials price and efficiency variances.
b. (Appendix) Prepare the journal entries to record the purchase and use of the direct materials using standard costing.
Answer #1

Actual quantities of direct materials used          30,800

 Actual costs of direct materials used           $137,800 

Standard price per unit of direct materials      $4.28 

 Flexible budget for direct materials              $121,300 


RA) 137,800/ 30,800= 4.474025974

(4.474025974- 4.28)* 30,800= 5,975.9999992 

or Answer: 5,976 U for DM price variance 


137,800- 121,300= 16,500

16,500- 5,976= 10,524 U for DM efficiency variance


RB) WIP inventory           DR             121,300

Materials price variance DR                5,976  from RA

Materials efficiency variance DR        10,524= 137,800- 121,300- 5,976

    Accounts payable   CR                                            137,800